# Exploiting Domain-Specific Parallel Data on Multilingual Language Models for Low-resource Language Translation

### v1

@@ A.3 Evaluation Metric
A.4 Jenson-Shannon Divergence
@@
Following our previous work [10], each corpus was first tokenized (times and numbers were also converted into the tokens <TIME> and <NUMBER>, respectively) using the NLTK package,131313https://www.nltk.org/ and stopwords were removed. The resulting corpora were transformed into a (discrete) frequency distribution over all word tokens. The frequency distributions of each train and test set are then compared using the formula above.
@@ Appendix B Pre-training results
Appendix C Tabular Results for Individual Techniques
C.1 Result Tables for EN-XX
@@ Appendix D Graphical Results of Performance
D.1 Single-domain ITTL Performance Plots
D.2 Muti-domain FT Performance Plots
D.3 Multi ITTL Performance Plots

### v2

@@ A.3. Evaluation Metric
A.4. Jenson-Shannon Divergence
@@
Following our previous work (Nayak et al., 2023), each corpus was first tokenized (times and numbers were also converted into the tokens <TIME> and <NUMBER>, respectively) using the NLTK package,141414https://www.nltk.org/ and stopwords were removed. The resulting corpora were transformed into a (discrete) frequency distribution over all word tokens. The frequency distributions of each train and test set are then compared using the formula above.
@@ Appendix B Pre-training results
Appendix C Tabular Results for Fine-tuning Techniques
| Fine-tuning Set | Fine-tuning Set Size | Telugu ID | Tamil ID | ||||
| FLORES | Bible | PMO | FLORES | Bible | PMO | ||
| PMO | 1k | 1.0 | 0.1 | 6.5 | 1.0 | 0.0 | 6.3 |
| 25k | 14.1 | 1.7 | 33.6 | 14.1 | 1.8 | 33.8 | |
| Bible | 1k | 0.3 | 8.2 | 0.2 | 0.2 | 7.8 | 0.2 |
| 25k | 2.2 | 27.8 | 0.9 | 2.1 | 27.8 | 0.9 | |
C.1. Result Tables for EN-XX
@@ Appendix D Graphical Results of Performance
D.1. Single-domain ITTL Performance Plots
Figure 4 and Figure 5 show the single-domain ITTL results for in-domain and out-domain cases (respectively).
D.2. Muti-domain FT Performance Plots
Figure 6 and Figure 7 show results of multi-domain FT for in-domain and out-domain cases (respectively).
D.3. Multi ITTL Performance Plots
Figure 8 and Figure 9 show results of multi-domain FT for in-domain and out-domain cases (respectively).